The average weight of 25 oranges selected from a normally distributed population is 400g with a standard deviation of 20g. Find the confidence interval using 95% confidence level.


1
Expert's answer
2022-04-29T09:18:15-0400

The critical value for "\alpha = 0.05" and "df = n-1 = 24" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.063899."

The corresponding confidence interval is computed as shown below:


"CI=(\\bar{X}-t_c\\times\\dfrac{s}{\\sqrt{n}},\t\\bar{X}+t_c\\times\\dfrac{s}{\\sqrt{n}})"

"=(400-2.063899\times\dfrac{20}{\sqrt{25}},"

"400+2.063899\times\dfrac{20}{\sqrt{25}})"

"=(391.7444,408.2556)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "391.7444<\mu<408.2556," which indicates that we are 95% confident that the true population mean "\u03bc" is contained by the interval "(391.7444,408.2556)."